I need to solve the following Linear Equation
-y=2x+1
-4x-2y=3

Answer by mananth(16946)

2 x + -1 y = 1 .............1
-1 y = 1 + -2 x
/ -1 =
y = -1 + 2 x
-4 x + -2 y = 3 .....................2
Substitute y in (2)
-4 x + -2 ( -1 + 2 x
-4 x + 2 + -4 x = 3
-4 x -4 x = 1
-8 x = 1
/ -8
x= - 1/ 8
Plug the value of x in (1)
2 x + -1 y = 1
2 * - 1/ 8 + -1 y = 1
- 1/ 4 + -1 y = 1
-1 y = 5/ 4
/ -1
y= -5/ 4
x=-1/8, y = -5/4